OpenStudy (anonymous):

A rectangular corral of widths Lx = L and Ly = 2.08L contains seven electrons. What multiple of h2/8mL2 gives the energy of the ground state of this system? Assume that the electrons do not interact with one another, and do not neglect spin.

OpenStudy (dominusscholae):

Energy of trap: \[E = nh^2(Lx^2+Ly2)/8mL^2\] =\[h^2(L^2+(2.08L)^2)/8mL^2 = 4.3264nh^2L^2/8mL^2 = 4.3264nh^2/8m\] Notice how we get an answer that is 4.3264 times h^2/8mL^2. Thus, 4.3264 would be your answer.

OpenStudy (dominusscholae):

^sorry 4.3264L

OpenStudy (dominusscholae):

^DOUBLE SORRY 4.3264L^2.

OpenStudy (anonymous):

the right answer is 16.7...

OpenStudy (dominusscholae):

:/ dang.